fill in the blank make the expression a perfect square v^2 - 6v +?
Answer by lwsshak3(11628)

fill in the blank make the expression a perfect square v^2 - 6v +?
take half the coefficient of the v term and square it.
add it to the expression then subtract the same amount to keep the expression in balance
(v^2 - 6v +9)-9=(v-3)^2-9
if the sign of the middle term is positive,
(v^2 + 6v +9)-9=(v+3)^2-9